solve f - 3 1/3 = 12 1/4
Answer by stanbon(75887)

solve f - 3 1/3 = 12 1/4
------------
f - (10/3) = (49/4)
---
Multiply thru by 12:
----
12f - 40 = 3*49
---
12f = 187
---
f = 15 29/50
=================
Cheers,
Stan H.